# DecoderTracker: Decoder-Only End-To-End method for Multiple-Object Tracking

### v5

@@ Abstract
Decoder-only methods, such as GPT, have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer methods. Over the years, end-to-end methods based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking. However,The substantial computational resource consumption of these methods, coupled with the optimization challenges posed by dynamic data, results in less favorable inference speeds and training times. To address the aforementioned issues, this paper optimized the network architecture and proposed an effective training strategy to mitigate the problem of prolonged training times, thereby developing DecoderTrack, a novel end-to-end tracking method. Subsequently, to tackle the optimization challenges arising from dynamic data, this paper introduced DecoderTrack+ by incorporating a Fixed-Size Query Memory and refining certain attention layers. Our methods, without any bells and whistles, outperforms MOTR on multiple benchmarks, with inference speeds 2.06 and 3.03 times faster than MOTR, respectively.
@@ 1 Introduction
Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images. The success of transformer [1] methods in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT methods like MOTR [2] have been introduced in the MOT domain. These methods offer advantages over the widely used tracking-by-detection methods [3, 4, 5, 6], such as eliminating the need for manual design or selection of features. Additionally, some enhanced methods [7, 8, 9] of MOTR have demonstrated the potential of this method class in terms of tracking accuracy.
@@
Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios. The number of detect queries is typically fixed, while the number of track queries fluctuates throughout the tracking process. This discrepancy results in dynamic data during inference, which introduces optimization challenges that limit the actual runtime speed of the method. To address this, we propose the Fixed-Size Query Memory (FSQM), which pads the tracking query sequence with empty queries to maintain a consistent number of queries during inference. By eliminating dynamic data, we have developed DecoderTracker+, which has significantly improved inference speed compared to MOTR and DecoderTracker.
At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the method to engage in less stringent learning might also be effective. Is there a possibility that this process could yield unexpected benefits? It is important to note that weak supervision, or self-supervised learning, has achieved significant success in many computer vision tasks [12, 13]. Based on this idea, we proposed the Tracking Box Selection Process (TBSP) which is mainly utilized during the training. Thus, DecoderTracker has been constructed by us.
After completing the aforementioned architectural optimizations mentioned above, we observed that while the training duration was significantly optimized, the inference speed did not improve as expected. This phenomenon prompted us to delve deeper into the issue. In theory, streamlining the network architecture should more notably enhance inference speed rather than training efficiency. To investigate this anomaly, we systematically analyzed the time consumption distribution across various modules. The results revealed that the decoder module accounted for the majority of the time, consuming nearly two-thirds of the model’s total inference time, which is evidently unreasonable. Through further cross-model comparative analysis, we identified that this issue also exists in MOTR but is absent in its base model, Deformable DETR. Controlled variable experiments indicated that fixing the number of queries significantly reduces decoder latency. Considering factors such as GPU memory usage, we ultimately pinpointed the root cause to be the dynamic query mechanism. This mechanism introduces unnecessary overhead during inference, such as inefficient GPU memory allocation in modern deep learning frameworks ,e.g., PyTorch/TensorFlow, and may also hinder compilation optimizations like operator fusion due to dynamic computation graphs. To address this issue, we proposed the Fixed-Size Query Memory (FSQM) for handling fixed-query tracking. By integrating FSQM and optimizing the attention layers in the original decoder and TAN modules, we introduced DecoderTrack+. Experimental results demonstrated that, although DecoderTrack+ theoretically increases computational load, its inference speed is significantly faster than the original version (from 19.6 fps to 28.8 fps) due to static data processing, while maintaining comparable tracking performance.
@@
    The second key innovation involves the introduction of a unique training strategy characterized by a three-stage process. At the same time, we incorporated TBSP (Temporal Balanced Sampling and Propagation), a weakly supervised training strategy designed for preliminary training of end-to-end MOT methods. These strategies accelerate method convergence, reduce overall training time, and enhance the efficiency of the training process.
@@
    Finally, by leveraging FSQM and optimizing certain self-attention layers within the network, we developed DecoderTrack+, which addresses the additional latency caused by dynamic data. This provides a fully viable technical pathway for the engineering deployment of MOTR-like end-to-end MOT models.
@@ 2.1 MOTR Series
Unlike tracking-by-detection methods exemplified by SORT [3, 14, 4], and previous transformer-based MOT methods [15, 16], MOTR [2] is the first method to achieve end-to-end multi-object tracking. It extends the object query from DETR [17] into a track query, thereby enabling tracking capabilities. It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing. Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for methoding long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration. These features establish MOTR as a significant advancement in multi-object tracking technology.
@@
Similarly, MeMOT [9] innovates by incorporating track queries that method and predict target trajectories iteratively across video frames. This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
MOTRv3 [8] and CO-MOT [20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework. MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance. In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces ”shadows” to enhance the diversity and quantity of queries, further refining the MOT capabilities.
@@ 2.2 Tracking by Detection
Tracking by Detection are among the most classical and widely used types of MOT. These methods typically employ a detector to first identify objects and then use various other methods for tracking.
In past years, the main methods within the Tracking by Detection category were SORT [3] and its derivative methods. The primary approach of these methods is to use a Kalman Filter to predict the positions of targets in the next frame and then perform data association with the detection results provided by the detector. Many subsequent improvements to these methods [5, 14, 4, 21] have introduced re-identification (reid) networks, incorporating object appearance features as an additional factor for association. There are also a series of methods like OC-SORT [22] that purely improve the Kalman Filter-based tracking method using kinematic methods.
In recent years, methods such as SUSHI [23], Glan [24] and PuTR [25] have emerged, which employ networks for tracking post-detection. The former two utilize Graph Neural Networks (GNN), while the last is based on transformer.
@@ 2.3 Real Time Detection Method
When discussing real-time detection methods, most people initially think of YOLO [26], which transformed the problem of object detection into a regression method, achieving efficient object detection. The method divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions. The most advanced versions in the YOLO series currently are YOLOv8.2 [27] and YOLOv9 [28].
Recently, the real-time detection method based on DETR, known as RT-DETR [29], has also demonstrated its extraordinary performance. RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the method structure.
Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary. Therefore, inspired by the YOLO series [28, 27], we designed a feature extraction network.
@@ 3.1 DecoderTracker
Figure 3: The structure of FENet. It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences. Specific module details can be found in Figure 4. The extracted features will be fed into the Decoder and introduced to the cross-attention module to interact with the queries. Similar to the Backbone and Encoder in the MOTR series, FENet focuses solely on extracting features from the current video frame, while the implicit spatio-temporal feature extraction is managed by the Decoder and the TAN.
3.1.1 DecoderTracker Architecture
DecoderTracker’s overall architecture is illustrated in the provided Figure 2. In this setup, a video sequence is input to image Feature Extraction Network(FENet) to extract features. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
For subsequent frames in the video sequence, the tracking queries from the previous frame and the learnable detect queries are concatenated and undergo filtering before being input to the decoder. These queries interact with the image features in the decoder to generate hidden states for bounding box predictions. Additionally, a feature aggregation process using MOTR’s TAN is performed to generate trajectory queries for the next frame.
@@ 3.1.2 Design of FENet
The structure of FENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction. The reason for not adopting the latest YOLOv9 or YOLOv10 is due to its slower training speed compared to YOLOv8. Through FENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
The design of this FENet is intended to enhance the method’s efficiency and increase its inference speed. Thus, our design primarily references the network structure of YOLOv8. The backbone structure of our method is almost identical to that of YOLOv8. Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
Figure 4: Specific module details, modules not specified in the diagram are network structures or APIs that come with PyTorch. These modules are primarily constructed based on CNN, capable of efficiently extracting image features. The employment of residual connections ensures that issues like gradient explosion and vanishing gradients are largely avoided.
@@ 3.1.5 Loss and Training Methods
$\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}+\lambda_{enc}\mathcal{L}_{enc}$ (3)
where $\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the frame predictions. The loss is calculated by ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$. $\mathcal{L}_{cls}$ denotes the focal loss[32], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[33]. $\mathcal{L}_{enc}$ is the anchor loss of DAB-DETR [31] and $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients, which are set to 1,2,5,5 in our method, just like DAB-DETR.
In this stage, the network is capable of initially learning the appearance features of objects to be tracked. During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time. Additionally, various data augmentation techniques [34, 35, 36] can be employed to expedite the method’s extraction of appearance features.
The second stage of training is conducted using MOTR’s Collective Average Loss. During this phase, only MOTR’s TAN and our TBSP are utilized. The loss functions used in this stage and the third stage are as Eq.4.
@@
At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects. Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage. Thanks to our various strategies and improvements to the network, we achieved huge improvements in training time and inference speed compared to MOTR.

### v6

@@ Abstract
Decoder-only methods, such as GPT, have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer methods. Over the years, end-to-end methods based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking. However,the substantial computational resource consumption of these methods, coupled with the optimization challenges posed by dynamic data, results in less favorable inference speeds and training times. To address the aforementioned issues, this paper optimized the network architecture and proposed an effective training strategy to mitigate the problem of prolonged training times, thereby developing DecoderTracker, a novel end-to-end tracking method. Subsequently, to tackle the optimization challenges arising from dynamic data, this paper introduced DecoderTracker+ by incorporating a Fixed-Size Query Memory and refining certain attention layers. Our methods, without any bells and whistles, outperforms MOTR on multiple benchmarks, featuring a 2 to 3 times faster inference than MOTR, respectively. The proposed method is implemented in open-source code, accessible at https://github.com/liaopan-lp/MO-YOLO.
@@ 1 Introduction
Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images. The success of transformer [1] methods in various domains in recent years is well documented, and transformer-based end-to-end MOT methods such as MOTR [2] and MOTIP [3] have recently been introduced in the MOT domain. These methods offer advantages over the widely used tracking-by-detecting (TBD) methods [4, 5, 6, 7], such as eliminating the need for manual design or feature selection. Among them, the MOTR series based on the Tracking-by-Query(TBQ) paradigm extends object queries to track queries and achieves tracking by propagating these track queries across frames. This category of methods exhibits exceptional extensibility, with their core ideas frequently adopted by models in multiple domains such as 3D detection[8], motion prediction[9], and video understanding[10]. In contrast, MOTIP[3], another end-to-end MOT method currently achieving higher accuracy, is somewhat inferior in this aspect. Its tracking mechanism is more akin to the TBD approach, except that its detection and tracking networks are jointly optimized during training. Thus, in-depth research on TBQ methods will not only significantly advance the development of MOT technology but also hold important implications for driving MOT toward greater efficiency and universality.
Figure 1: Illustration of diverse MOT pipelines. For MOTIP and TBD, a detector first yields detections, followed by ID assignment networks or matching associators for tracking. In contrast, MOTR-series tracking-by-query methods use track queries to model targets and propagate them frame-by-frame. Our approach also employs tracking-by-query but features a simpler network; the proposed FSQM overcomes dynamic data challenges inherent in other tracking-by-query methods.
However, the inference speed of these TBQ methods is often suboptimal. Their RNN-like linear inference structure makes them less GPU-friendly during training, and this linear training paradigm also renders data augmentation techniques commonly used in detection tasks inapplicable. Upon in-depth analysis, it becomes clear that the core mechanism of TBQ is accomplished by the decoder, modeling object trajectories and temporal correlations through query propagation and attention interactions, while the encoder merely handles static image feature extraction and contributes little to the core logic of dynamic tracking.
@@
At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the method to engage in less stringent learning might also be effective. Is there a possibility that this process could yield unexpected benefits? It is important to note that weak supervision, or self-supervised learning, has achieved significant success in many computer vision tasks [14, 15]. Based on this idea, we proposed the Tracking Box Selection Process (TBSP) which is mainly utilized during training. Thus, we propose DecoderTracker, a highly efficient end-to-end TBQ tracking network. Its fundamental departure from prior approaches is visually discernible in Figure 1.
After completing the aforementioned architectural optimizations mentioned above, we observed that while the training duration was significantly optimized, the inference speed did not improve as expected. This phenomenon prompted us to delve deeper into the issue. In theory, streamlining the network architecture should more notably enhance inference speed rather than training efficiency. To investigate this anomaly, we systematically analyzed the time consumption distribution across various modules. The results revealed that the decoder module accounted for the majority of the time, consuming nearly two-thirds of the model’s total inference time, which is evidently unreasonable. Through further cross-model comparative analysis, we identified that this issue also exists in MOTR but is absent in its base model, Deformable DETR. Controlled variable experiments indicated that fixing the number of queries significantly reduces decoder latency. Considering factors such as GPU memory usage, we ultimately pinpointed the root cause to be the dynamic query mechanism. This mechanism introduces unnecessary overhead during inference, such as inefficient GPU memory allocation in modern deep learning frameworks ,e.g., PyTorch/TensorFlow, and may also hinder compilation optimizations like operator fusion due to dynamic computation graphs. To address this issue, we proposed the Fixed-Size Query Memory (FSQM) for handling fixed-query tracking. By integrating FSQM and optimizing the attention layers in the original decoder and TAN modules, we introduced DecoderTracker+. Experimental results demonstrated that, although DecoderTracker+ theoretically increases computational load, its inference speed is significantly faster than the original version (from 19.6 fps to 28.8 fps) due to static data processing, while maintaining comparable tracking performance.
@@
    The second key innovation involves the introduction of a unique training strategy characterized by a three-stage process. At the same time, we incorporated TBSP, a weakly supervised training strategy designed for preliminary training of end-to-end MOT methods. These strategies accelerate method convergence, reduce overall training time by enhancing the efficiency of the training process.
@@
    Finally, by leveraging FSQM and optimizing certain self-attention layers within the network, we developed DecoderTracker+, which addresses the additional latency caused by dynamic data. This provides a fully viable technical pathway for the engineering deployment of MOTR-like end-to-end MOT models.
@@ 2.1 MOTR Series
Unlike tracking-by-detection methods exemplified by SORT [4, 16], and previous transformer-based MOT methods [17, 18], MOTR [2] is the first method to achieve end-to-end multi-object tracking. It extends the object query from DETR [19] into a track query, thereby enabling tracking capabilities. It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing. Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for methoding long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration. These features establish MOTR as a significant advancement in multi-object tracking technology.
@@
Similarly, MeMOTR [22] innovates by incorporating track queries that predict target trajectories iteratively across video frames. This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
MOTRv3 [23] and CO-MOT [24] focus on optimizing query interactions and label assignment within the MOTR framework. MOTRv3 adopts a release-fetch supervision strategy for balanced label assignments, while CO-MOT employs cooperative label assignment with ”shadows” to enhance query diversity. DNMOT[25] addresses severe occlusions through denoising training (simulating occlusion with noise) and a Cascaded Mask Module that coordinates query interactions to prevent suppression of occluded trajectories. DQFormer[26], meanwhile, uses Decoupled Query Augmentation to separate detection and association query branches, reducing feature conflicts. A
@@ 2.2 Tracking by Detection
Tracking by Detection frameworks are among the most classical and widely used approaches for MOT. These methods typically employ a detector to first identify objects and then use various other methods for tracking.
In past years, the main methods within the TBD category were SORT [4] and its derivative methods. The primary approach of these methods is to use a Kalman Filter to predict the positions of targets in the next frame and then perform data association with the detection results provided by the detector. Many subsequent improvements to these methods [6, 16, 7] have introduced re-identification (reid) networks, incorporating object appearance features as an additional factor for association. There are also a series of methods like OC-SORT [27] and MotionTrack[5] that purely improve the Kalman Filter-based tracking method using kinematic methods. In recent years, within the TBD paradigm, a new wave of methods has emerged, such as SportMamba[28], TWiX[29], MSPNet[30] and Glan[31]. These approaches leverage advanced neural network architectures—including spatiotemporal modeling networks and query-based attention mechanisms—to refine the post-detection tracking process, moving beyond the traditional reliance on Kalman filtering or standalone re-identification modules. Unlike earlier methods that primarily relied on kinematic predictions or appearance features for association, these newer frameworks integrate deeper spatiotemporal reasoning into the tracking pipeline, enabling more robust handling of complex scenarios such as long-term occlusion and fast motion.
@@ 2.3 Real Time Detection Method
When discussing real-time detection methods, most people initially think of YOLO [32], which transformed the problem of object detection into a regression method, achieving efficient object detection. The method divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions. Among recent variants in the YOLO family, YOLOv8 and its upgraded version YOLO11[33] stand out for their superior stability and migration capability, making them the most reliable choices for practical deployment.
In recent years, a real-time detection method based on DETR, known as RT-DETR [34], has also shown extraordinary performance. RT-DETR has also demonstrated extraordinary performance, achieving optimized inference speeds by designing an efficient encoder and simplifying its network architecture. In particular, YOLO11 primarily focuses on modifications in the head section, while its core feature extraction framework remains largely consistent with YOLOv8. Given the proven stability and mature migration performance of YOLOv8 in various scenarios, our work thus draws more on YOLOv8 for FENet design, ensuring reliability in our MOT pipeline, which prioritizes low latency over the transformer encoder.
@@ 3.1 DecoderTracker
3.1.1 Architecture of Proposed Methods
The overall architecture of DecoderTracker and DecoderTracker+ are illustrated in the provided Figure 2. Each frame of the input video is fed into the Feature Extraction Network(FENet) to obtain multi-scale feature maps. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is inputted to the decoder. For subsequent frames in the video sequence, the tracking queries from the previous frame and the learnable detect queries are concatenated and undergo filtering before being input to the decoder. These queries interact with the image features in the decoder to generate hidden states for bounding box predictions. Additionally, MOTR’s Track Association Network (TAN) is employed to update the track queries, providing trajectory queries for the next frame.
@@ 3.1.2 Design of FENet
Figure 3: The structure of FENet. It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences. The left part of the figure corresponds to the backbone, which consists of C2m and Conv3 modules (detailed in Figure 4) to extract initial multi-scale features from input images. The right part represents the neck, where features from the backbone are processed through upsampling and concatenation operations—specifically, high-resolution features from upper layers are fused with low-resolution features from lower layers via C2m modules—to generate the final three-scale feature maps. These feature maps are then fed into the decoder and introduced to the cross-attention module to interact with the queries. Similar to the Backbone and Encoder in the MOTR series, FENet focuses solely on extracting features from the current video frame, while the implicit spatio-temporal feature extraction is managed by the Decoder.
FENet is a lightweight backbone-plus-FPN network derived from YOLOv8. It replaces the ResNet-50 backbone and the multi-layer transformer encoder in MOTR, delivering three-scale feature maps to the decoder. The reason for not adopting the latest YOLOv9 or YOLOv10 is due to its slower training speed compared to YOLOv8. Through FENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder along with the track and detect queries.
Our FENet is intended to enhance the efficiency of the method and increase its inference speed. Thus, our design primarily resembles the YOLOv8 network structure. The only differences are in the neck structure, which we adapt to fit the input requirements of the decoder.
Figure 4: Specific module details. Modules not specified in the diagram are network structures or APIs that come with PyTorch. The C2m module, composed of Split, BottleNeckF, and Concat operations, is a core component in both the backbone and neck of FENet (as shown in Figure 3), responsible for feature refinement and propagation. The SPPF module, located in the backbone, uses multiple MaxPool2d layers to enhance multi-scale feature aggregation. These modules are primarily constructed based on CNN, capable of efficiently extracting image features. The employment of residual connections ensures that issues like gradient explosion and vanishing gradients are largely avoided.
@@ 3.1.5 Loss and Training Methods
$\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\text{giou}}\mathcal{L}_{\text{giou}}+\lambda_{\text{enc}}\mathcal{L}_{\text{enc}}$ (3)
where $\widehat{Y}=\{\widehat{Y}_{i}\}_{i=1}^{N}$ denotes frame predictions, $Y=\{Y_{i}\}_{i=1}^{N}$ is the ground truth, and $\omega=\{\omega_{i}\}_{i=1}^{N}$ represents matched indices. The component losses are defined as follows:
$\mathcal{L}_{\text{cls}}$ $=\sum_{i=1}^{N}-\alpha\left(1-p_{i,c_{i}}\right)^{\gamma}\log\left(p_{i,c_{i}}\right)\quad$ (4)
[…]
$\mathcal{L}_{\text{giou}}$ $=\sum_{i=1}^{N}\left[1-\text{GIoU}\left(\hat{b}_{i},b_{i}\right)\right]\quad$ (6)
$\mathcal{L}_{\text{enc}}$ $=\sum_{j=1}^{K}\left[\left\|\hat{b}^{\text{enc}}_{j}-b_{\sigma(j)}\right\|_{1}+\beta\left(1-\text{GIoU}\left(\hat{b}^{\text{enc}}_{j},b_{\sigma(j)}\right)\right)\right]\quad$ (7)
Here, $p_{i,c_{i}}$ denotes the predicted probability for the ground truth class $c_{i}$ at the matched index $i$. The predicted and ground truth bounding boxes are represented by $\hat{b}_{i}$ and $b_{i}$, respectively, each defined by the center coordinates, width, and height. The $j$-th encoder-predicted anchor box is denoted as $\hat{b}^{\text{enc}}_{j}$. The matching process is governed by $\sigma$, a function that maps the top-$K$ encoder predictions to ground truth instances based on center distance, where $K=\min(\text{top-}K,N)$ is dynamically set to the number of ground truth objects. Hyperparameters include $\alpha=0.25$ and $\gamma=2$ for focal loss and $\beta=1$ for GIoU weight in anchor loss. The loss weights are configured as $\lambda_{\text{cls}}=1$, $\lambda_{l_{1}}=2$, $\lambda_{\text{giou}}=5$, and $\lambda_{\text{enc}}=5$ to balance classification, L1 regression, GIoU, and encoder losses. In addition, a more specific explanation of these losses can be found in [38, 39, 37].
In this stage, the network is able to initially learn the appearance characteristics of the objects to be tracked. During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time. Furthermore, various data augmentation techniques [40, 41, 42] can be used to expedite the method’s extraction of appearance features. Whether these data augmentation methods are applied has a significant impact on the results of this stage, exerting an approximate one point effect on the final tracking accuracy. Since this is not the focus of this paper, detailed results are not enumerated here.
The second stage of training is carried out using the Collective Average Loss of MOTR. During this phase, only MOTR’s TAN and our TBSP are utilized. The loss functions used in this stage and the third stage are as in Eq.8.
@@
At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects. Newly appeared objects in the second stage are defined as high-confidence predictions from detect queries (exceeding a confidence threshold) that pass the TBSP filtering process, i.e., their bounding boxes do not significantly overlap (IOU $\leq$ $\delta_{iou}$) with any existing track query (see Section 3.1.3 for details on TBSP). This ensures that they are distinct from already tracked objects. Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to that of the second stage. Thanks to our various strategies and improvements to the network, we achieved huge improvements in training time and inference speed compared to MOTR.